# thanks: These authors contributed equally to this workthanks: These authors contributed equally to this work

### v1

@@ I Introduction
The spin polarization of quasiparticles in solid-state materials is a key property for many quantum physical phenomena, not the least in the field of spintronics. Conventional antiferromagnetic materials feature spin-degenerate electron bands, as dictated by parity-time-reversal ($\mathcal{PT}$) symmetry. However, it has recently been understood that antiferromagnetic materials can display a momentum-dependent spin-splitting which is distinct from relativistically spin-orbit coupled systems. One such mechanism, originally envisioned in Ref. [1], consists of a spin-lattice coupling due to an internal periodic magnetic field in the material. When $\mathcal{PT}$ symmetry is broken by the combined lattice geometry and spin ordering in a material, a large spin splitting can arise which depends on momentum $\boldsymbol{k}$ [2, 3, 4, 5], while maintaining zero net magnetization upon integration over the entire first Brillouin zone  [^1] According to the extended classification in Ref. [6], altermagnets with vanishing magnetization belong to types II and III altermagnets. Such materials have been named altermagnets [7], and ab initio calculations have identified several possible material candidates that can host an altermagnetic state. This includes metals like RuO2 [3, 8, 9] and Mn5Si3 [10] as well as semiconductors/insulators like MnTe [9, 7], CrSb [7], MnF2 [4, 11], and La2CuO4 [7]. Recent ARPES measurements in MnTe [12, 13, 14], RuO2 [15, 16], and CrSb [17, 18, 19] have corroborated several of these predictions  [^2] The magnetic order in RuO2 is debated [20, 21].. Another piece of evidence supporting altermagnetism is provided by measuring the anomalous Hall effect [22, 23] and spin-splitter effect and torques [24, 25, 26]. It is noteworthy that a Fermi-surface instability caused by quasiparticle interactions has been predicted to cause an altermagnetic spin-splitting in the band structure [27] long before the recent surge of activity in the field. On the other hand, such an instability is interestingly restricted for $p$-wave magnetism [28].
@@ III.1 Tight-binding $s-d$ model for antiferromagnet and altermagnet
We have explicitly set $\boldsymbol{S}_{i}\propto\hat{\boldsymbol{z}}$. Moreover, we note that the antiferromagnetic state can be modelled with a simpler lattice, namely a 2-site unit cell where the normal atoms in our model have been omitted. We have confirmed that the essential physical mechanisms underlying the interface-induced magnetization, which will be explained below, are also present in such a model.
@@
The characterization of magnetization and edge effects revolves around the study of the spin-resolved electron density and how it varies across the lattice. The eigenfunctions of our lattice, $\psi_{n,k,\alpha,\sigma}(r_{I})$, are obtained from the column vectors diagonalizing $H$ and are resolved at the unit cell level and each eigenfunction is described by three unit cell indices $k,\alpha,\sigma$ as well as an x-coordinate I. The spin- and $\alpha$-dependent density in the unit cell at position $r_{I}$ is defined as
$n_{I,\alpha,\sigma}=\sum_{n,k}|\psi_{n,k,\alpha,\sigma}({r}_{I})|^{2}n_{F}(\varepsilon_{n,k,\alpha,\sigma}),$ (35)
where $\varepsilon_{n,k,\alpha,\sigma}$ is the energy of the eigenfunction $\psi_{n,k,\alpha,\sigma}$.
@@
Figure 4: NM densities obtained by a sum over the two lowest energy harmonics for $k=0$ shown in the unit cell-resolved (a) and site-resolved (b) case. Analogous plots in (c) and (d) are shown in the case of a non-zero on-site coupling $\varepsilon_{2}=\delta$, $\varepsilon_{3}=-\delta$ for $\delta\ll t$. (a) In the NM, the eigenfunctions are pair-wise equivalent with the ones corresponding to sites in the left column (blue) and right column (red) equal. The left and right column eigenfunctions are shifted by the lattice constant $a$, corresponding to their relative displacement within the unit cell. Note that the $a$ shift between columns is present in all eigenfunctions and therefore carries over to the density. (b) The site-resolved density can be illustrated by adding the densities for $\alpha=1$ and $\alpha=2$ for the row 1 and $\alpha=3$ and $\alpha=4$ for the 2. row. (c) A change in the on-site potentials changes the magnitude of the eigenfunctions, lifting the previous equivalence between the densities at different sites $\alpha$. The probability amplitude at $\alpha=2$ sites is suppressed due to $\varepsilon_{2}=\delta$, while the negative potential $\varepsilon_{3}=-\delta$ enhances the probability amplitude at the $\alpha=3$ sites. The site-resolved densities are given by the densities of unit cell sites within the same row, shifted relatively by a lattice constant $a$. Due to the induced magnitude difference, the system density becomes a $\pi$-periodic density wave.
@@
Before moving on, we briefly mention that in the following, we will use the terminology unit cell-resolved and site-resolved extensively. Site-resolved refers to quantities that are functions of individual lattice site indices $(i,k)$ while unit cell-resolved refers to quantities which are functions of unit cell indices $(I,\alpha,k)$.
@@
Figure 5: Density profiles for the split system $\varepsilon_{2(3)}=\pm\delta$ showing Friedel oscillations at $\mu=-1.5t$. Both the unit cell-resolved density (a) and site-resolved density (d) show clear Friedel oscillations, but the site-resolved density also show the $\pi$-density wave due to the shift in eigenfunction magnitudes. The Fourier transform of the unit cell-resolved density is shown in (b) and the site-resolved in (e). Several peaks arise in both corresponding to the dominant frequencies in the density oscillations. We label the dominant peaks with the notation $2k_{F,i}$ where i is a simple index. Due to the 2D geometry, we do not have one particular, but rather a distribution of $k_{F}^{x}$ from states from different y-modes. We quantify this distribution by summing the row-resolved bulk spectral function evaluated at the Fermi level across all y-momenta, shown in (f) where $k_{x}$ runs over the full Brillouin zone (FBZ), obtaining a $2k_{F}^{x}$ distribution which matches well with the observed frequency content in the density oscillations. Note that values of $2k_{F}^{x}$ beyond $\pi/2$ appear in the reduced Brillouin zone (RBZ) in (c) due to aliasing and is distinguished by an asterix ($2k_{F}^{*}$). (e) In the site-resolved density, an additional set of frequencies given by $|2k_{F}\pm\pi|$ arise due to the fact that the staggered $\pi$ order is modulated in magnitude by the underlying eigenfunctions.
@@
Consider now a normal metal (NM) lattice with $J_{\text{sd}}=0$. We consider two distinct configurations of the non-magnetic on-site potentials in order to elucidate the interplay between the unit cell potentials and the lattice geometry. In Fig. 4, we plot the unit cell-resolved and site-resolved densities for a system with no on-site potentials in (a) and (b) and a split potential $\varepsilon_{2(3)}=\pm\delta$ in (c) and (d) where $\delta/t\ll 1$ (see the insets for unit cell layout). The densities are summed over the two first standing-wave harmonics in x for the homogeneous y-mode ($k=0$). The first observation is that the eigenfunctions corresponding to sites in the left and right column of the unit cell are shifted relatively by a lattice constant $a$ [see Fig. 4(a) and (c)] due to the relative x-displacement between sites in the two columns, causing the magnitudes of the eigenfunctions to be different. In a unit cell picture where both columns correspond to the same indices $(I,k)$, this magnitude difference manifests as a lattice constant $a$ shift between the eigenfunctions. The second observation is that the magnitudes of the individual eigenfunctions are altered by the respective on-site potentials. In the case of $\varepsilon_{2}=\varepsilon_{3}=0$, all sites in the unit cell are equivalent causing the only distinguishing factor to be the one lattice constant $a$ shift between eigenfunctions in different columns. For the split configuration ($\varepsilon_{2(3)}=\pm\delta$), this equivalence is lifted. In Fig. 4(c), the eigenfunction corresponding to $\alpha=2$ is suppressed as $\varepsilon_{2}=\delta$ constitute an effective lowering of the chemical potential. Along the same lines, an enhancement is observed for $\alpha=3$ due to $\varepsilon_{3}=-\delta$. The magnitudes of the $\alpha=1,4$ eigenfunctions also change due to a changed unit cell environment, but the effect is equal across the two sites and comparatively smaller. To understand the significance of these relative shifts in magnitude, the site-resolved picture is illustrative. The site-resolved densities are obtained by plotting the densities of unit cell sites in the same row ($\alpha=1,2$ and $\alpha=3,4$) together, shifted relatively by a lattice constant $a$, shown in Figs. 4(b) and 4(d). As is evident from Fig. 4(b), the unit cell description of the NM with $\varepsilon_{2}=\varepsilon_{3}=0$ does not essentially provide any new information not present in the original site-resolved model. The two unit cell-resolved densities come together to form the equally shaped density profile obtained in the original site-resolved model. The advantage of the unit-cell picture arise when on-site potentials are introduced. The periodic modulation caused by the split potentials in Figs. 4(c) and (d) give rise to a $\pi$-periodic density wave on each row in the unit cell. When the sum in Eq. (35) is extended to a large number of eigenfunctions, the site-resolved density profiles can become intricate. As we have seen, however, the observed oscillations can be easily deconvoluted in the unit cell picture as the combination of a one lattice constant $a$ shift and different eigenfunction magnitudes, caused by the chosen on-site potential configuration and unit cell geometry.
@@
In a system with open boundaries, the oscillations of the charge density known as Friedel oscillations [52, 53] arise in the vicinity of the interfaces. While the oscillations of eigenfunctions interfere and cancel out in the bulk of the ribbon, the open boundaries force all eigenfunctions to be roughly in phase close to the boundaries leading to the Friedel oscillations. The frequency of 1D Friedel oscillations is $2k_{F}$ where $k_{F}$ is the Fermi momentum. In 2D, the Fermi surface does in general give rise to a distribution of Fermi wave vectors, and the whole distribution of $k_{F}^{x}$, the x component of the Fermi wave vectors, can leave a mark on the density oscillations.
@@
To exemplify how these density oscillations occur, the density for the split potential ($\varepsilon_{2(3)}=\pm\delta$) is shown in Fig. 5 for the unit-cell- 5(a) and site-resolved 5(d) representations. Both show clear oscillations in the vicinity of the interface while roughly homogeneous in the ribbon bulk. In Figs. 5(b) and 5(e), the Fourier transform of the density in the two representations is shown, indicating the dominant frequencies in the density oscillations. These frequencies correspond to roughly twice the x-component of the Fermi vector for the homogeneous y-mode. Due to the 2D Fermi surface, however, eigenfunctions with a range of $k_{F}^{x}$ exist at the Fermi surface, giving rise to additional peaks in the frequency spectrum. These can be shown to correspond closely to the distribution of $k_{F}^{x}$ in the partially integrated spectral function evaluated at the Fermi surface, obtained by integrating over all y-momenta, show in Figs. 5(c) and 5(f). Finally, we point out that the density oscillations arising from the Fermi surface are largely independent on the specific on-site potentials as long as these are small relative to t. As long as the Fermi surface does not change significantly upon changing the potentials, the system eigenfunctions will exhibit similar Friedel oscillations. The non-zero potentials leave a more direct mark on the site-resolved density oscillations [Fig. 5(e)] through the emergence of a $\pi$-peak, due to the previously discussed mechanism, and through the appearance of additional peaks at frequencies $|2k_{F}\pm\pi|$, caused by the $2k_{F}$ modulation of the staggered order. The staggered order is established by the difference in magnitude between eigenfunctions which will oscillate with the same frequency as that of the underlying eigenfunctions, $2k_{F}$. This gives rise to the additional peaks in the Fourier transform of the site-resolved density in Fig. 5(d) close to $k_{x}=0$.
@@
The purpose of this section is to identify the key mechanisms at play in the microscopic model concerning density modulations, which are summarized in the paragraph above. With an understanding of the mechanisms that determine the oscillation pattern of the electron density near an interface, we are now ready to move on to discuss the consequences of spin-dependent potentials in the unit cell.
Figure 6: Spin-up and spin-down densities are shown in (a) for an AFM with $J_{\text{sd}}=t$, $\mu=-3.5t$. The spin-dependent $J_{\text{sd}}$ gives rise to highly spin-dependent occupations on the magnetic sites $\alpha=1,4$. (b) The magnitude differences in the underlying eigenfunctions gives rise to a staggered magnetization (green and orange). The magnetization is out of phase between the two rows causing the total unit cell magnetization (black) to approach zero in the ribbon center. Despite no spin-dependence in the Friedel oscillations, we still observe uncompensated magnetization (black) in the vicinity of the interface. Note that we here plot the magnitude of the unit cell magnetization multiplied by a factor 5 for clarity. This is caused by an effective coupling between the staggered order and the spin-independent Friedel oscillations, not captured in the continuum model.

### v2

@@ I Introduction
The spin polarization of quasiparticles in solid-state materials is a key property for many quantum physical phenomena, not the least in the field of spintronics. Conventional antiferromagnetic materials feature spin-degenerate electron bands, as dictated by parity-time-reversal ($\mathcal{PT}$) symmetry. However, it has recently been understood that antiferromagnetic materials can display a momentum-dependent spin-splitting which is distinct from relativistically spin-orbit coupled systems. One such mechanism, originally envisioned in Ref. [1], consists of a spin-lattice coupling due to an internal periodic magnetic field in the material. When $\mathcal{PT}$ symmetry is broken by the combined lattice geometry and spin ordering in a material, a large spin splitting can arise which depends on momentum $\boldsymbol{k}$ [2, 3, 4, 5, 6, 7], while maintaining zero net magnetization upon integration over the entire first Brillouin zone  [^1] According to the extended classification in Ref. [8], altermagnets with vanishing magnetization belong to types II and III altermagnets. Such materials have been named altermagnets [9, 10], and ab initio calculations have identified several possible material candidates that can host an altermagnetic state [11]. This includes metals like RuO2 [3, 7, 12] and Mn5Si3 [13] as well as semiconductors/insulators like MnTe [12, 9], CrSb [9], MnF2 [4, 14], and La2CuO4 [9]. Recent ARPES measurements in MnTe [15, 16, 17], RuO2 [18, 19], and CrSb [20, 21, 22] have corroborated several of these predictions  [^2] The magnetic order in RuO2 is debated [23, 24].. Another piece of evidence supporting altermagnetism is provided by measuring the anomalous Hall effect [25, 26] and spin-splitter effect and torques [27, 28, 29]. It is noteworthy that a Fermi-surface instability caused by quasiparticle interactions has been predicted to cause an altermagnetic spin-splitting in the band structure [30] long before the recent surge of activity in the field. On the other hand, such an instability is interestingly restricted for $p$-wave magnetism [31].
@@ III.1 Tight-binding $s-d$ model for antiferromagnet and altermagnet
Figure 4: The spectral function $A_{\sigma}(k_{x},E)$ is shown for $k_{y}=0$ in the model with translational invariance in both x and y. The NM band structure is shown in (a) where the energy states between $E\sim-2t$ and $E\sim 2t$ are backfolded into the reduced Brillouin zone (RBZ) due to the unit cell segmentation. Introducing a non-zero $J_{\text{sd}}=t$, gaps open at the AFM zone edge (b). If one additionally breaks the equivalence between the two non-magnetic sites, a spin-split altermagnetic band structure emerges (c), in this case by setting $\varepsilon_{3}=100t$. The insets show the Fermi surface at the chemical potential denoted by the dotted line.
We have explicitly set $\boldsymbol{S}_{i}\propto\hat{\boldsymbol{z}}$. Moreover, we note that the antiferromagnetic state can be modelled with a simpler lattice, namely a 2-site unit cell where the normal atoms in our model have been omitted. We have confirmed that the essential physical mechanisms underlying the interface-induced magnetization, which will be explained below, are also present in such a model. To elucidate the spectral information of this model, the spectral functions are plotted for a bulk model (translational invariant in both x and y) for $k_{x}$ vs $E$ in Fig. 4 showing the effect of a non-zero $J_{\text{sd}}$ and a broken equivalence between $\varepsilon_{2}$ and $\varepsilon_{3}$ on the spectral properties of the model.
@@
The characterization of magnetization and edge effects revolves around the study of the spin-resolved electron density and how it varies across the lattice. The eigenfunctions of our lattice, $\psi_{n,k,\sigma}(r_{I},\alpha)$, are obtained from the column vectors of the matrix $U$ diagonalizing the Hamiltonian, and are resolved at the unit cell level. Each eigenfunction is described by three indices $k,n,\sigma$ and is a function of the x-coordinate I and the site index $\alpha$ within each unit cell. The spin- and $\alpha$-dependent density at x-position $r_{I}$ in the ribbon is defined as
$n_{I,\alpha,\sigma}=\sum_{n,k}|\psi_{n,k,\sigma}({r}_{I},\alpha)|^{2}n_{F}(\varepsilon_{n,k,\sigma}),$ (35)
where $\varepsilon_{n,k,\sigma}$ is the energy eigenvalue of the eigenfunction $\psi_{n,k,\sigma}$.
@@
Figure 5: NM densities obtained by a sum over the two lowest energy harmonics for $k=0$ shown in the unit cell-resolved (a) and row-resolved (b) case. Analogous plots in (c) and (d) are shown in the case of a non-zero on-site coupling $\varepsilon_{2}=\delta$, $\varepsilon_{3}=-\delta$ for $\delta\ll t$. (a) In the NM, the densities are pair-wise equivalent with the ones corresponding to sites in the left column (blue) and right column (red) equal. The left and right column eigenfunctions are shifted by the lattice constant $a$, corresponding to their relative displacement within the unit cell. (b) The row-resolved density can be illustrated by adding the densities for $\alpha=1$ and $\alpha=2$ for the row 1 and $\alpha=3$ and $\alpha=4$ for the 2. row. (c) A change in the on-site potentials changes the magnitude of the eigenfunction entries corresponding to the particular site, lifting the previous equivalence between the densities at different sites $\alpha$. The probability amplitude at $\alpha=2$ sites is suppressed due to $\varepsilon_{2}=\delta$, while the negative potential $\varepsilon_{3}=-\delta$ enhances the probability amplitude at the $\alpha=3$ sites. The row-resolved densities are given by the densities of unit cell sites within the same row, shifted relatively by a lattice constant $a$. Due to the induced magnitude difference, the system density becomes a $\pi$-periodic density wave.
@@
Before moving on, we briefly mention that in the following, we will use the terminology unit cell-resolved and row-resolved extensively. Row-resolved refers to quantities that are functions of individual lattice site indices $(i,k)$ across one of the two distinct rows on our ribbon (first and second row in the unit cell, see Fig. 1), while unit cell-resolved refers to quantities which are functions of unit cell indices $(I,\alpha,k)$.
@@
Figure 6: Density profiles for the split system $\varepsilon_{2(3)}=\pm\delta$ showing Friedel oscillations at $\mu=-1.5t$. Both the unit cell-resolved density (a) and row-resolved density (d) show clear Friedel oscillations, but the row-resolved density also show the $\pi$-density wave due to the presence of non-zero on-site potentials. The Fourier transform of the unit cell-resolved density is shown in (b) and the row-resolved in (e). Several peaks arise in both corresponding to the dominant frequencies in the density oscillations. We label the dominant peaks with the notation $2k_{F,i}$ where i is a simple index. Due to the 2D geometry, we do not have one particular, but rather a distribution of $2k_{F}^{x}$ from states from different y-modes. We quantify this distribution by summing the row-resolved bulk spectral function evaluated at the Fermi level across all y-momenta, shown in (f) where $k_{x}$ runs over the full Brillouin zone (FBZ), obtaining a $2k_{F}^{x}$ distribution which matches well with the observed frequency content in the density oscillations. Note that values of $2k_{F}^{x}$ beyond $\pi/2$ appear in the reduced Brillouin zone (RBZ) in (c) due to aliasing and is distinguished by an asterisk ($2k_{F}^{*}$). (e) In the row-resolved density, an additional set of frequencies given by $|2k_{F}\pm\pi|$ arise due to the fact that the staggered $\pi$ order is modulated in magnitude by the underlying eigenfunctions.
@@
Consider now a normal metal (NM) lattice with $J_{\text{sd}}=0$. We consider two distinct configurations of the non-magnetic on-site potentials in order to elucidate the interplay between the unit cell potentials and the lattice geometry. In Fig. 5, we plot the unit cell-resolved and row-resolved densities for a system with no on-site potentials in (a) and (b) and a split potential $\varepsilon_{2(3)}=\pm\delta$ in (c) and (d) where $\delta/t\ll 1$ (see the insets for unit cell layout). The densities are summed over the two first standing-wave harmonics in x for the homogeneous y-mode ($k=0$). The first observation is that the densities on sites in the left and right column of the unit cell are shifted by a lattice constant $a$ [see Fig. 5(a) and (c)]. In a unit cell picture where both columns correspond to the same indices $(I,k)$, this shift originating from the relative x-displacement between columns, manifests as a magnitude difference between the density profiles. The second observation is that the magnitudes of the $\alpha$-resolved densities are altered individually by the respective on-site potentials. In the case of $\varepsilon_{2}=\varepsilon_{3}=0$, all sites in the unit cell are equivalent causing the only distinguishing factor to be the one lattice constant $a$ shift between densities in different columns. For the split configuration ($\varepsilon_{2(3)}=\pm\delta$), this equivalence is lifted. In Fig. 5(c), the $\alpha=2$ density is suppressed as $\varepsilon_{2}=\delta$ constitute an effective lowering of the chemical potential. Along the same lines, an enhancement is observed for $\alpha=3$ due to $\varepsilon_{3}=-\delta$. The magnitudes of densities at sites $\alpha=1,4$ sites also change due to a changed unit cell environment, but the effect is equal across the two sites and comparatively smaller. To understand the significance of these relative shifts in magnitude, the row-resolved picture is illustrative. The row-resolved densities are obtained by plotting the densities of unit cell sites in the same row ($\alpha=1,2$ and $\alpha=3,4$) together, shifted relatively by a lattice constant $a$, shown in Figs. 5(b) and 5(d). As is evident from Fig. 5(b), the unit cell description of the NM with $\varepsilon_{2}=\varepsilon_{3}=0$ does not essentially provide any new information not present in the original model. The advantage of the unit-cell picture arise when on-site potentials are introduced. While the unit cell-resolved densities (Fig. 5(c)) are modulated in magnitude due to the on-site potentials, they retain their smooth standing wave-like profiles as all sites of the same $\alpha$ see the same potential change. The row-resolved density, however, is given by alternating left and right column sites, and the row-resolved densities in Fig. 5(d) are now distinctly different, where a $\pi$-periodic density wave emerge on each row due to the alternating magnitude modulation. When the sum in Eq. (35) is extended to a large number of eigenfunctions, the row-resolved density profiles can become intricate. As we have seen, however, the observed oscillations can be deconvoluted in the unit cell picture as the combination of a one lattice constant $a$ shift and magnitude differences between densities in different columns, caused by the chosen on-site potential configuration and unit cell geometry.
@@
In a system with broken translation invariance the oscillations of the charge density known as Friedel oscillations [55, 56] arise in the vicinity of interfaces and defects. While the oscillations of eigenfunctions interfere and cancel out in the bulk of the ribbon, the open boundaries force all eigenfunctions to be roughly in phase close to the boundaries leading to the Friedel oscillations. The frequency of 1D Friedel oscillations is $2k_{F}$ where $k_{F}$ is the Fermi momentum. In 2D, the Fermi surface does in general give rise to a distribution of Fermi wave vectors, and the whole distribution of $k_{F}^{x}$, the x component of the Fermi wave vectors, can leave a mark on the density oscillations.
@@
To exemplify how these density oscillations occur, the density for the split potential ($\varepsilon_{2(3)}=\pm\delta$) is shown in Fig. 6 for the unit-cell- 6(a) and row-resolved 6(d) representations. Both show clear oscillations in the vicinity of the interface while roughly homogeneous in the ribbon bulk. In Figs. 6(b) and 6(e), the Fourier transform of the density in the two representations is shown, indicating the dominant frequencies in the density oscillations. These frequencies correspond to roughly twice the x-component of the Fermi vector for the homogeneous y-mode. Due to the 2D Fermi surface, however, eigenfunctions with a range of $k_{F}^{x}$ exist at the Fermi surface, giving rise to additional peaks in the frequency spectrum. These can be shown to correspond closely to the distribution of $k_{F}^{x}$ in the partially integrated spectral function evaluated at the Fermi surface, obtained by integrating over all y-momenta, show in Figs. 6(c) and 6(f). Note in particular that the presence of the split potentials break the symmetry between the rows. As such, the row-resolved Fermi-surfaces are identical in shape, but with a slight difference in the distribution of states on the Fermi surface. This is evident in Fig. 6(c) and 6(f) [see insets in 6(f)] where the distributions of x-components are different on the two rows.
Finally, we point out that the density oscillations arising from the Fermi surface are largely independent on the specific on-site potentials as long as these are small relative to t. As long as the Fermi surface does not change significantly upon changing the potentials, the system eigenfunctions will exhibit similar Friedel oscillations, but with an additional staggered order on top. The non-zero potentials leave a more direct mark on the row-resolved density oscillations [Fig. 6(e)] through the emergence of a $\pi$-peak, due to the previously discussed mechanism, and through the appearance of additional peaks at frequencies $|2k_{F}\pm\pi|$, caused by the $2k_{F}$ modulation of the $\pi$ density wave magnitude. In effect, the presence of periodic potentials on the lattice gives rise to a magnitude modulation on top of the standing wave eigenfunctions. The magnitude of these modulations follow the magnitude of the eigenfunction as a whole, causing a coupling between the underlying $2k_{F}$ frequency of the initial eigenfunction, and the $\pi$-frequency arising from the lattice geometry. This gives rise to the additional peaks in the Fourier transform of the row-resolved density in Fig. 6(d) close to $k_{x}=0$.
@@
The purpose of this section is to identify the key mechanisms at play in the microscopic model concerning density modulations. With an understanding of the mechanisms that determine the oscillation pattern of the electron density near an interface, we are now ready to move on to discuss the consequences of spin-dependent potentials in the unit cell.